HUBBLE SPACE TELESCOPE DAILY REPORT #845 PERIOD COVERED: 0800 EST Daily Status Report as of 076/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 3791 (Quasar Absorption Line Survey) The Faint Object Spectrograph was used to make five observations of quasar PG1216+069 to find out if this is an appropriate target for future observations. The observations were executed in course track. There was a reasonable amount of signal in the first four observations. However, the peakup did not straddle the target, so it was not cretin that the target was well position in the slit. The last observations received a number of communication data hits. The spectrum looked good, though it is not certain that the target was well centered in the aperture. 1.2 Completed WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make internal flats observations. These observations will be used for the restoration of Earth flats, which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. The observations were reported to look okay. Some of the telemetry communications were reported to be garbled. 1.3 Completed PC 4714 (WF/PC Calibration: PC Internal Delta Flats) The Planetary Camera was used to make seven internal flat observations. All seven of the observations were received and displayed in the OSS. The observations was reported to look okay. However, some of the data segments had been garbled in transmission. 1.4 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat calibration images which will be used to look for the return of quantum efficiency hysteresis and new forms of contamination (i.e., measles and daisies). Both sets of internal flat observations were reported to look normal. 1.5 Completed PC 4347 (WF/PC Calibration: Flood Retention Test) The Planetary Camera was used with filter F439W to take three Earth flat calibration images. One of the flats was reported to be underexposed, and the other two were saturated. 1.6 Completed GHRS/2 1168 (Interstellar Carbon and Oxygen) The High Resolution Spectrograph was used to attempt observations of standard star HD37043 to obtain high signal to noise ratio data which will hopefully show narrow interstellar absorption lines of carbon and oxygen. The Baseline guide star acquisition failed. and the first onboard target acquisition did not occur (See Operations Note 2.4). An attitude reference update was sent to the spacecraft, and the peakup target acquisition was successful. It is not certain that the star was well centered in the aperture for the spectral observations. 1.7 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.8 Completed PC 4140 (WF/PC Calibration: Rapid Internal Monitor, Part 2) The Planetary Camera was used to make internal flat observations using filter F349W to look for form of quantum efficiency hysteresis. The observations were executed as scheduled and no problems were reported. 1.9 Completed GHRS/2 3496 (Ultraviolet Atlas of Sirius) The High Resolution Spectrograph (WSCAN Mode) was used to obtain ultraviolet spectra of the star Sirius (aka, HD48915) to get actuate flux distribution and line profiles. The observations were executed as scheduled, and no problems were reported. 1.10 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies wit FOC/96) The Faint Object Camera (f/96) was used to make ultraviolet snapshot of the nearby galaxy UGC-1655. The observations were monitored by the OSS, and assessed the data quality. No problems were reported. 1.11 Initiated WFC 4658 (Determination of Extragalactic Distance Scale) Observations have been initiated using the Wide Field Camera to look at two different fields within galaxy M-101 (aka, NGC-457) to help locate Cepheid variable stars. The observations are currently ongoing, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 2 (CT at 076/0213Z, See Operations Note 2.4) Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 12 Successful : 7 (Failures at 076/0110, 0114,0128, 0148, and 0151 See Operations Note 2.4) 2.3 Significant HSTARs: None 2.4 Operations Note: One Guide Star acquisition and five Fixed Head Star Tracker (FHST) updates failed due to attitude error introduced during the HIGH MODE DATA TAKE interval. An HSTAR (3929) was written, and the problem is being investigated. The subsequent reacquisition was saved by an Attitude Reference Update Planned Real-Time Slew (ARU/PRT). 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4658 (Determination of Extragalactic Distance Scale) 3.2 Complete FOC/96/WFC 3325 (Structure of Globular Clusters) 3.3 Initiate GHRS/2 3755 (Absorption Cross-section of Nearby Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: FGS #2 Anomaly Investigation Further testing and analysis of the FGS #2 anomalous behavior are still ongoing. In the interim FGS #2 will not be used for observations. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-